# StableQ: Enhancing Data-Scarce Quantization with Text-to-Image Data

### 07 Dec 2023 v1

@@ 4.2 Few-shot StableQ
[…]
in the style of $\texttt{in the style of \{}\mathtt{S}\texttt{\}"}.$" . (9)
@@
where $M$ is the number of all object classes. This method is similar to the Textual Inversion [21] which uses a similar technique to learn the token of an object across several images. In contrast to [21], our method learns the token across multiple objects, aiming to characterize the whole dataset. Also, considering that we only have 1 image in each class, performing class-wise Textual Inversion may suffer from over-fitting and generate less diverse data (Fig. 2\⃝raisebox{-0.9pt}{3}).
Parameterized BatchNorm Distribution Filtering. The previous BN Sensitivity filter in Zero-shot StableQ does not require any data. To further enhance the quality of generated data, we leverage the few-shot training data to train a parameterized BN distribution filter (Fig. 2\⃝raisebox{-0.9pt}{4}). As demonstrated in [4], channels in deep neural networks are correlated and of different importance. Hence, instead of simply averaging all BN channels and layers together (cf. Eq. (missing) 7), we train a filter model $g(\cdot)$ that takes the BN distance of each layer and outputs whether the image should be kept or removed.
To train this parameterized filter, we leverage the few training data as in-distribution samples and also craft out-of-distribution samples by randomly permuting the pixels of the training data. As a result, we optimize the filter to output 1 or 0, which corresponds to in-distribution samples (kept for quantization) or out-of-distribution samples (removed).
The filter is lightweight, using only a simple MLP structure, and can be trained with few input samples to significantly improve the filtering quality. In addition, the inference can be done with image-by-image prediction, i.e., batch size is 1. The overall StableQ pipeline is described in Fig. 2 for both zero-shot and few-shot image synthesis.
Figure 3: Visualization of zero/few-shot StableQ and real images of high-speed train on ImageNet dataset.
Table 1: Evaluation of zero-shot post-training quantization on CNN models (top-1 accuracy (%)).
Quant. Method Syn. Method #Bits (W/A) ResNet-18 ResNet-50 MobileNetV2 MobileNet-b MnasNet-1.0
@@
    The figures are taken from  [39]. ${}^{\dagger}$ It synthesizes $5K$ images while others synthesize only $1K$ images.
    Denotes our implementation based on the open-source code, hyper-parameters are kept the same for all data synthesis methods.
4.3 StableQ Finetuning
To further improve the performance of PTQ models, QAT-based finetuning will be benefecial. In this work, we take the state-of-the-art PTQ pipeline [38, 48, 36] and apply QAT finetuning with StableQ generated synthetic data. Specifically, after rounding learning in PTQ, the quantization becomes
where $\mathrm{sgn}(\cdot)$ is the sign function indicating rounding up or down. The vector ${\bm{v}}$ is already well-optimized in the PTQ reconstruction stage. To initialize from the PTQ model and stabilize the training process, we detach all previous learnable variables including ${\bm{w}}$ and ${\bm{v}}$, and reinitialize an all-zero vector ${\bm{u}}$. The new quantization function is thus given by
The de-quantization step remains the same with Eq. (missing) 2. By introducing an additional variable ${\bm{u}}$ and detaching the previous variable, we start from the original PTQ model and avoid passing gradient through floor operation. As a result, the ${\bm{u}}$ can be safely updated through STE.
5 Experiments
In this section, we empirically demonstrate the superiority of our StableQ synthetic data in both a qualitative and a quantitative way. For the PTQ scenario, we follow the conventional setup [38, 36, 39] to synthesize 1k images. As for the QAT scenarios, we synthesize 1.2M images with 1200 images in each object class to match the original ImageNet dataset. Unless specified, we use Stable Diffusion v1-5 [1] and set the guidance scale to 3.5222StableQ generates one image per second on one A100 GPU.. We will first provide the visualization of our synthetic data, and then compare them against existing state-of-the-art methods. Finally, we analyze StableQ by conducting various ablation and generalization studies.
5.1 Analysis on Synthetic Data
As shown in Fig. 1, the StableQ images after filtering are high-quality and visually similar to real images. Moreover, to show the effect of our few-shot image generation, we provide an example of high-speed train class in the ImageNet dataset [13]. Fig. 3 shows the 1-shot real image, the zero-shot label prompt images, and the few-shot token prompt inversion images. The zero-shot images have more natural sites as background, while the few-shot images generate more in-station views just like the real image.
Table 2: Evaluation of zero-shot post-training quantization on ViT models (top-1 accuracy (%)).
Quant. Method Syn. Method #Bits ViT-B DeiT-B Swin-B
Full Prec. N/A 32/32 84.54 81.80 85.27
PTQ4ViT [69] Real Data 4/4 58.07 63.57 75.20
PSAQ-ViT [40] 60.26 65.18 72.12
StableQ 63.17 67.08 74.27
@@ 5.1 Analysis on Synthetic Data
PSAQ-ViT [40] 60.18 74.69 54.21
StableQ 67.50 76.10 70.08
Table 3: Evaluation of zero-shot quantization-aware training CNNs (top-1 accuracy (%)).
Method #Real Data #Syn Data #Bits (W/A) ResNet-18 MobileNetV2 ResNet-50
LSQ [18] (ICLR 2020) 1.2M 0 4/4 71.10 69.50 76.70
@@
ZAQ [43] (CVPR 2021) 0 4.6M 52.64 0.10 53.02
Qimera [11] (NeurIPS 2021) 0 1.2M 63.84 61.62 66.25
IntraQ [75] (CVPR 2022) 0 5k 66.47 65.10 -
ARC+AIT [12] (CVPR 2022) 0 1.2M 65.73 66.47 68.27
DSG [53] (TPAMI 2023) 0 1.2M 62.18 60.46 71.96
AdaDFQ [52] (CVPR 2023) 0 1.2M 66.53 65.41 68.38
TexQ [9] (NeurIPS 2023) 0 1.2M 67.73 67.07 70.72
StableQ 0 1.2M 70.03 69.65 76.10
LSQ [18] (ICLR 2020) 1.2M 0 3/3 70.20 65.30 75.80
@@
AdaDFQ [52] (CVPR 2023) 0 1.2M 38.10 28.99 17.63
TexQ [9] (NeurIPS 2023) 0 1.2M 50.28 32.80 25.27
StableQ 0 1.2M 68.18 59.15 73.99
5.2 Zero-Shot PTQ Evaluation
We evaluate our proposed method by testing it on two types of neural networks, (1) ConvNets such as ResNet [23], MobileNet [30, 58], RegNet [55], and MnasNet [61] and (2) Vision Transformers like original ViT [16]. For CNNs PTQ, we employ the state-of-the-art PTQ algorithms—Brecq [38] and Genie-M [36] to perform the W4A4 and W2A4 quantization, which reconstruct the activation output in a block-wise manner. In this case, we use 1024 synthetic images for quantization evaluation. For ViT PTQ, we perform PTQ4ViT [69] and RepQ-ViT [41] across 5 different models, including ViT-M [16], DeiT-M [63], and Swin-B [44]. We follow the open-source code implementation and generate 32 images as the calibration dataset to obtain the quantized ViTs. For all cases, we generate 2$\times$ synthetic data and filter 50% of them to match the volume.
Comparison on CNNs. We select existing state-of-the-art image synthesis algorithms as our comparison, including ZeroQ [8], the Knowledge Within [25], IntraQ [75], Qimera [11], MixMix [39], Genie-D [36]. For reference, we also include the performance of PTQ using 1024 real ImageNet-1k images. We summarize all the accuracy results in Table 1. It can be observed that our StableQ achieves the highest accuracy in most cases. Interestingly, we find that the network architecture affects the performance of synthetic data. On ResNet architectures, StableQ and Genie-D have similar accuracy (0.1-0.5%) levels, however, StableQ largely outperforms other zero-shot algorithms on lightweight network architectures due to their lower quantization resilience. As an example, StableQ increases 5% accuracy compared to Genie-D on MobileNet-b W2A4 quantization and 4.4% accuracy on MNasNet W2A4 quantization as well.
Comparison on Vision Transformers. Given that there is only one zero-shot algorithm for ViTs, PSAQ-ViT [40], we compare StableQ with PSAQ-ViT in Table 2. StableQ consistently outperforms PSAQ-ViT [40]. For example, on ViT-B and Swin-B, our method has 7.3% and 15% accuracy improvement using the RepQ-ViT method. This result proves that StableQ can achieve state-of-the-art performance on both CNNs and ViTs.
Table 4: Evaluation of few-shot quantized ResNet-18 (W3A3).
Method #Real Data #Syn Data Accuracy
Genie-M [36] 1k 0 67.34*
[…]
10k 0 67.94*
100k 0 68.32*
300k 0 69.13*
1.2M 0 70.20
[…]
    denotes our implementation.
5.3 Zero-shot QAT Evaluation
@@
We then compare StableQ with existing generative zero-shot QAT methods, such as GDFQ [65], ZAQ [43], Qimera [11], IntraQ [75], ARC [12], AdaDFQ [52], TexQ [9]. Note that these methods jointly optimize GAN and quantized model. Thus, they can generate unlimited synthetic data during finetuning. We originally have 1.6M synthetic images (1600 images/class), and then filter 0.4M (400 images/class) images using the energy filter and the BN distance filter. We use the SGD optimizer with a learning rate of 0.001 followed by a cosine annealing decay schedule for 50 epochs. Additionally, we report the LSQ results using 1.2M real ImageNet data. The results are summarized in Table 3. Our method largely outperforms other zero-shot QAT methods and nearly approaches the accuracy of LSQ baseline. Remarkably, StableQ improves the accuracy of W3A3 quantization by a large margin, for instance, StableQ exceeds 48% accuracy of TexQ, the best-performing method.
5.4 Few-shot QAT Evaluation
@@
In this section, we evaluate the few-shot QAT scenarios. Our main comparison is LSQ [18] using different amounts of real data. We show how much of the real data can our StableQ generated synthetic data match in practice. To ensure a fair comparison, we initialize the QAT model by performing Genie-M on 1k real data and then finetune the model using {1k, 10k, 100k, 300k, 1.2M} real data, corresponding to 1-shot, 10-shot, 100-shot, 300-shot, and full-shot data regimes, respectively. To generate StableQ data, we initialize the token embedding as ImageNet, and train it for 50k iterations with a batch size of 8, learning rate set to 5e-4 with Adam optimizer.
As summarized in Table 4, the initialized PTQ model reaches 67.34% accuracy, and the few-shot LSQ increases its accuracy as more real images are provided. QAT with only 1k real images does not bring much improvement (0.1%). The performance of QAT is highly correlated with the number of real data. Finally, despite only using 1-shot real data, StableQ is able to achieve approximately 100-shot QAT accuracy.
@@ 5.5 Data Transferability Evaluation
Table 5: Cross-model evaluation on synthetic data. We compare the data transferability between existing (a) and our (b) methods.
(a) Genie-D & PSAQ-ViT
Source model Res18 MBV2
Res18 64.86 48.37
MBV2 64.21 51.47
ViT-B-16 47.27 17.96
(b) StableQ
Source model Res18 MBV2
Res18 65.72 54.32
MBV2 65.52 54.82
ViT-B-16 65.44 54.24
[39] has demonstrated that the synthetic data extracted from one model has low transferability on other models, i.e., low data transferability among different models. To demonstrate that StableQ has relatively high transferability, we conduct a cross-model evaluation, e.g., synthesizing images based on model A, and test the quantization on model B. In 5(a) we provide the results of Genie-D and PSAQ-ViT, tested on zero-shot PTQ W2A4 scenarios. We then summarize the results of StableQ in 5(b). We observe that the StableQ data, although filtered by a different model, only drops 0.2-0.5% final accuracy, while the existing methods drop 0.6-33% accuracy, especially when using ViT synthetic data for MobileNetV2 quantization.
@@ 5.6 Ablation study
Number of synthetic data in PTQ. In practice, synthesizing images using StableQ does not cost a lot of computing resources (1 image/second). Hence, we can safely increase the number of synthetic data with minimal overhead in PTQ cases. Yet retrieving more real data seems much more difficult if the training data is private or under IP protection. In Fig. 4, we show that increasing the synthetic data can consistently improve the PTQ performances of W2A4 quantized ResNet-50. StableQ even outperforms real-data baseline when using 4k input images.
Figure 4: Accuracy vs # synthetic data with zero-shot PTQ.
Figure 5: Accuracy vs filtering strength with zero-shot PTQ.
Energy Score Filtering Strength. Given a fixed number of synthetic data ($N$) used for quantization and a percentage $r$ representing how many images are filtered out, the total amount of generated images is $\frac{N}{1-r}$. We test different options of $r$ from $\{0.1,0.3,0.5,0.7,0.9\}$ to test the effect of energy score filtering. We choose ResNet-18 and use W2A4 zero-shot PTQ for evaluation. The test performance is shown in Fig. 5(A). We generally find that a higher filtering ratio leads to better test accuracy. However, it will also increase the total number of generated images. Nevertheless, the accuracy variation is rather small.
BN Sensitivity Filtering Strength. We run the same test for the BN sensitivity filtering mechanism. The results are shown in Fig. 5(B). Unlike the energy score, setting the ratio to 50% has the best performance, we hypothesize that the networks need more diverse data as stated in [72].

### 08 Mar 2024 v2

@@ 4.2 Data-Scarce GenQ
$\mathrm{Prompt}_{[i]}=\texttt{"A photo of a }\{\mathtt{C}_{[i]}\}\ \texttt{in the style of \{}\mathtt{S}\texttt{\}"}.$" . (10)
@@
where $M$ is the number of all object classes. This method learns the token across multiple objects, aiming to characterize the whole dataset [^2] We provide more technical details in Appendix A. . Fig. 2\⃝raisebox{-0.9pt}{3} illustrates the data-scarce prompt optimization. After optimization, we directly used the learned token embedding and the class label to generate images and filter the output images using the same technique in Fig. 2\⃝raisebox{-0.9pt}{2}.
4.3 Quantization with Synthetic Data from GenQ
The selected synthetic data from GenQ is then used to quantize a full-precision pre-trained model using QAT or PTQ. For PTQ, we adopt the state-of-the-art reconstruction-based rounding optimization methods [34, 32]. As for QAT, we propose to finetune the quantized network on top of a PTQ model. Specifically, after the rounding optimization in PTQ, the quantization becomes
where $\mathrm{sgn}({\bm{v}})$ is the learned rounding indicating up or down, which is already well-optimized in the PTQ reconstruction stage. To initialize from the PTQ model and stabilize the training process, we freeze all previous learnable variables including ${\bm{w}}$ and ${\bm{v}}$, and reinitialize an all-zero vector ${\bm{u}}$. The new quantization function is thus given by
The de-quantization step remains the same with Eq. (missing) 2. By introducing an additional variable ${\bm{u}}$ and freezing the previous variable, we start from the original PTQ model and avoid passing the gradient through floor operation. As a result, the ${\bm{u}}$ can be safely updated through STE in finetuning. Note, during PTQ or QAT, we only use the synthetic data from GenQ to perform quantization.
5 Experiments
In this section, we empirically demonstrate the effectiveness and the efficiency of our GenQ synthetic data in both a qualitative and a quantitative way. For the PTQ scenario, we follow the conventional setup [34, 32, 35] to synthesize 1k images. As for the QAT scenarios, we synthesize 1.2M images with 1200 images in each object class to match the original ImageNet dataset volume. Unless specified, we use Stable Diffusion v1-5 [1] and set the guidance scale to 3.5. We will first provide the visualization of our synthetic data, the latency comparison of generating the data, and then compare them against existing state-of-the-art methods across multiple setups. Finally, we analyze GenQ by conducting various ablation studies. All experiments and accuracy noted are for the ImageNet dataset.
Figure 3: Visualization of DF/DS GenQ and existing data synthesis method for quantization in low data regimes.
5.1 Analysis on Synthetic Data
Visualization. As shown in Fig. 3, the GenQ images after filtering are high-quality and visually similar to real images. To compare the synthetic data with existing data-free quantization, we also provide the images from [50, 35, 32]. It can be observed that our method can generate much higher quality synthetic data than previous inversion-based methods. For example, [50] only demonstrates certain textures rather than objects, [35] shows some objects but the scene shows less resolution and clarity.
Data Synthesis Cost. We further study the cost of each synthetic data generation method. Specifically, we measure the latency of Genie [32], TexQ [8], and our method corresponding to image inversion, GAN-based, and text-to-image synthesis approaches, respectively, for 1000 images on the ImageNet dataset. The pre-trained model used for filtering or input optimization is ResNet-50. We report the latency in 1(a), from which we can find using text-to-image synthesis provides faster data generation than image inversion-based techniques. Although GAN-based data generation is faster than GenQ , their data is not reusable across model types, which increases the overall latency if many models have to be quantized.
Table 1: Comparison of synthetic data generation approaches w.r.t. (a) generation speed and (b) data quality.
(a) Latency
Method Type Latency Reusable
Genie [32] Image inversion 152 min
TexQ [8] GAN-based 10 min
GenQ Text-to-image 24 min
(b) BatchNorm Distance
Data Source Res50 MBV2
Genie [32] 0.014$\pm$0.001 1.03$\pm$0.296
TexQ [8] 0.045$\pm$0.013 N/A
GenQ 0.024$\pm$0.002 0.78$\pm$0.106
Data Quality Assessment. We further measure the quality of the synthetic data through some quantitative evaluation metrics. Specifically, we measure the average BN distances of the generated data as well as the original training data on ResNet-50 and MobileNetV2. The results are shown in 1(b), from which we can find that GenQ largely closes the gap between synthetic data and real training data.
Table 2: Evaluation of data-free PTQ on CNN models (top-1 accuracy (%)).
Quant. Method Syn. Method #Bits W/A Res18 Res50 MBV2 MB-b MNas-1
@@ 5.1 Analysis on Synthetic Data
    The figures are taken from  [35].
    Denotes our implementation based on the open-source code.
Table 3: Evaluation of data-free QAT on CNN models (top-1 accuracy (%)).
Method #Real Data #Syn Data #Bits W/A Res18 MBV2 Res50
LSQ [15] 1.2M 0 4/4 71.10 69.50 76.70
@@
ZAQ [39] 0 4.6M 52.64 0.10 53.02
Qimera [10] 0 1.2M 63.84 61.62 66.25
IntraQ [72] 0 5k 66.47 65.10 -
ARC+AIT [11] 0 1.2M 65.73 66.47 68.27
DSG [50] 0 1.2M 62.18 60.46 71.96
AdaDFQ [49] 0 1.2M 66.53 65.41 68.38
TexQ [8] 0 1.2M 67.73 67.07 70.72
GenQ 0 1.2M 70.03 69.65 76.10
LSQ [15] 1.2M 0 3/3 70.20 65.30 75.80
@@
AdaDFQ [49] 0 1.2M 38.10 28.99 17.63
TexQ [8] 0 1.2M 50.28 32.80 25.27
GenQ 0 1.2M 68.18 59.15 73.99
5.2 Quantization Performance Evaluation
In this section, we test the quantization accuracy with our synthetic data and compare it with various existing methods. We will introduce the experiment setup in each scenario individually.
Comparisons on Data-Free PTQ. We start with evaluating our proposed method by testing it on data-free PTQ including (1) CNNs such as ResNet [19], MobileNet [26, 54], and MnasNet [57] and (2) Vision Transformers like original ViT [14]. For PTQ on CNNs, we employ the state-of-the-art PTQ algorithms, Brecq [34] and Genie-M [32] to perform the W4A4 and W2A4 quantization, which reconstruct the activation output in a block-wise manner. In this case, we use 1024 synthetic images for quantization evaluation. For PTQ on ViT, we perform PTQ4ViT [66] and RepQ-ViT [37] across 5 different models, including ViT-M [14], DeiT-M [59], and Swin-B [40]. We follow the open-source code implementation and generate 32 images as the calibration dataset to obtain the quantized ViTs. For all cases, we generate 2$\times$ synthetic data and filter 50% of them to match the volume.
For CNNs, we select SOTA existing methods including ZeroQ [7], the Knowledge Within [21], IntraQ [72], Qimera [10], MixMix [35], Genie-D [32]. For reference, we also include the performance of PTQ using 1024 real ImageNet-1k images. We summarize all the accuracy results in Table 2. It can be observed that our GenQ achieves the highest accuracy in most cases. Interestingly, we find that the network architecture affects the performance of synthetic data. On ResNet architectures, GenQ and Genie-D have similar accuracy (0.1-0.5%) levels, however, GenQ largely outperforms other data-free algorithms on lightweight network architectures due to their lower quantization resilience. As an example, GenQ increases 5% accuracy compared to Genie-D on MobileNet-b W2A4 quantization and 4.4% accuracy on MnasNet W2A4 quantization.
Given that there is only one data-free algorithm for ViTs, PSAQ-ViT [36], we compare GenQ with PSAQ-ViT in Table 4. GenQ consistently outperforms PSAQ-ViT [36]. For example, on ViT-B and Swin-B, our method has 7.3% and 15% accuracy improvement using the RepQ-ViT method. This result proves that GenQ can achieve state-of-the-art performance on both CNNs and ViTs.
@@ 5.2 Quantization Performance Evaluation
Comparisons on Data-Free QAT. We then compare GenQ with existing data-free QAT methods, such as GDFQ [62], ZAQ [39], Qimera [10], IntraQ [72], ARC [11], AdaDFQ [49], TexQ [8]. Note that these methods jointly optimize GAN and quantized model. Thus, they can generate unlimited synthetic data during finetuning. We originally have 1.6M synthetic images (1600 images/class), and then filter 0.4M (400 images/class) images using energy score and BN distance filtering. We use the SGD optimizer with a learning rate of 0.001 followed by a cosine annealing decay schedule for 50 epochs of QAT. Additionally, we report the LSQ results using 1.2M real ImageNet data. The results are summarized in Table 3. Our method largely outperforms other data-free QAT methods and nearly approaches the accuracy of LSQ baseline. Remarkably, GenQ improves the accuracy of W3A3 quantization by a large margin, for instance, GenQ exceeds the accuracy of TexQ, the best-performing method, by 48% on ResNet-50.
@@
Comparisons on Data-Scarce QAT. In this section, we evaluate the data-scarce QAT scenarios. Our main comparison is LSQ [15] using different amounts of real data. We show how much of the real data can our GenQ generated synthetic data match in practice. We experiment this on a ResNet-50 with W3A3 quantization. We initialize the QAT model by performing Genie-M on 1k real data and then finetune the model using {1k, 10k, 100k, 300k, 1.2M} real data, corresponding to 1-shot, 10-shot, 100-shot, 300-shot, and full-shot data regimes, respectively. To generate GenQ data, we initialize the token embedding as ImageNet, and train the embedding for 50k iterations, (see Appendix A for more details).
As demonstrated in Fig. 4, the performance of QAT is highly correlated with the number of real data. Our method can boost the performance of QAT under the low data regime. For example, when only 1k real images are provided (i.e., 1-shot), GenQ can achieve 68.88% ImageNet accuracy, similar to the QAT performance that needs 300$\times$ more training data.
@@ 5.3 Data Transferability Evaluation
[Uncaptioned image] Figure 5: Accuracy vs # syn. data with data-free PTQ using Genie-M. Table 5: Cross-model evaluation on synthetic data. We compare the data transferability between (a) existing methods and (b) GenQ . (a) Genie-D & PSAQ-ViT Source model Res18 MBV2 Res18 64.86 48.37 MBV2 64.21 51.47 ViT-B-16 47.27 17.96 (b) GenQ Source model Res18 MBV2 Res18 65.72 54.32 MBV2 65.52 54.82 ViT-B-16 65.44 54.24
It is demonstrated that the synthetic data extracted from one model has low transferability on other models [35]. To demonstrate that GenQ has relatively high transferability, we conduct a cross-model evaluation, e.g., synthesizing images based on model A, and evaluate the quantization on model B. In Table 5(a) we provide the results of Genie-D and PSAQ-ViT, tested on data-free PTQ W2A4 scenarios. We then summarize the results of GenQ in Table 5(b). We use Genie-M as our PTQ method for CNN architectures and PASQ-ViT for ViT. We observe that the GenQ data, although filtered by a different model, only drops 0.2-0.5% final accuracy, while the existing methods drop 0.6-33% accuracy, especially when using ViT synthetic data for MobileNetV2 (MBV2) quantization.
@@ 5.4 Ablation study
Number of synthetic data in PTQ. In practice, synthesizing images using GenQ requires very low latency (1 image/second). Hence, we can safely increase the number of synthetic data with minimal overhead in PTQ cases. Yet retrieving more real data seems much more difficult if the training data is private or under IP protection. In Fig. 5, we show that increasing the synthetic data can consistently improve the PTQ performances of W2A4 quantized ResNet-50. GenQ even outperforms the quantized model baseline (optimized using PTQ on the 1k real training data) when using only 4k synthetic input images.
Figure 6: Accuracy vs filtering strength with data-free PTQ.
Energy Score Filtering Strength. Given a fixed number of synthetic data ($N$) used for quantization and a percentage $r$ representing how many images are filtered out, the total amount of generated images is $\frac{N}{1-r}$. We test different options of $r$ from $\{0.1,0.3,0.5,0.7,0.9\}$ to test the effect of energy score filtering. We choose ResNet-18 and use W2A4 data-free PTQ (with Genie-M) for evaluation. The test performance is shown in Fig. 6(a). We generally find that a higher filtering ratio leads to better test accuracy. However, it will also increase the number of generated images. Nevertheless, the accuracy variation is rather small.
BN Sensitivity Filtering Strength. We run the same test for the BN sensitivity filtering mechanism. The results are shown in Fig. 6(b). Unlike the energy score, setting the ratio to 50% has the best performance. We hypothesize that the networks need more diverse synthetic data to effectively quantize a model.